About 97% of Earth’s water is found in the oceans. Another fraction is present as water vapor

in the atmosphere (clouds). Some water is contained in the solid state as ice and snow in

snowpacks, glaciers, and the polar ice caps. Surface water is found in lakes, streams, and

reservoirs. Groundwater is located in aquifers underground. And minor fraction is stored in

dams or dikes. Like Tarbela dam is 9.679 million acre feet originally, has been decreased to

33.5%. Suppose water content having 145.6 ppm total dissolved solid, 47.8 mg/L is its Nacl

contents, however the sodium is only 15.0ppm, Potasium is 7.0 mg/L while total hardness is

71.0 mg/L. You are required to solve given below problem by using these concentrations.

Q-2a) Why Tarbela reservoir losses its capacity? Could you calculate if the same is

incremental decrease by a decade when this reservoir will completely dry up/no storage

capacity?

Q-2b) Calculate (what if Tarbela reservoir is over capacity) the total salt content. One thing

you may be remember that the capacity has been decrease 33.5%.

Q-2c) If the Tarbela Lake has 390 tons of Salt in tem of NaCl, KCl, and CaCl, how huge

quantity of water may be evaporated to acquire this quantity of salt?


1
Expert's answer
2021-02-01T03:44:13-0500

2a.) Tarbela dam losses it's capacity due to sedimentation for over a long period.

If,

"10yrs=3.242

 \newline ?? =6.437"

"\dfrac{6.437*10}{3.242}=19yrs"


2b).total salt content at 33.5%

After a decrease of 33.5% reminder is

"9.679=145.6ppm \newline 6.437=??"


"\dfrac{6.437*145.6}{9.679}"


"=96.83ppm"